OpenStudy (anonymous):

Find the missing monomial −6x^10.?=84^23

Directrix (directrix):

@shakur I do not know what this means: −6x^10.?=84^23 Would you consider clicking on the blue DRAW tab below and writing the problem? Thanks.

OpenStudy (anonymous):

\[-6^{10}.?=84^{23}\]

OpenStudy (anonymous):

Find the missing Monomial

OpenStudy (anonymous):

\[−6x ^{10}.?=84^{23}\]

OpenStudy (anonymous):

@Directrix did you get the question

Directrix (directrix):

Yes, I understand what the question is. Thanks.

OpenStudy (anonymous):

The Question find the missing monomial \[-6x ^{10}.?=84^{23}\]

hero (hero):

Okay, so have you isolated x yet?

OpenStudy (anonymous):

I havent

hero (hero):

In fact, I'm pretty sure this is not possible unless you know the value of x

OpenStudy (anonymous):

Oh I think I did mistake on the question let me retype it

OpenStudy (anonymous):

\[-6x^{10}.?=84x ^{23} \]

OpenStudy (anonymous):

Find the missing monomial

hero (hero):

Good luck with that one. I figured out what x equals.

OpenStudy (anonymous):

What about this one multiply 3(7x+6)

OpenStudy (anonymous):

Its also polynomial question

hero (hero):

This one is a lot easier

OpenStudy (anonymous):

Do help me with it ,do the steps for me and the answer

hero (hero):

a(b + c) = ab + ac

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

whats did you get

OpenStudy (anonymous):

3(7x+6)= 21x+18

OpenStudy (anonymous):

Am I right or wrong

hero (hero):

It's right.

hero (hero):

Good luck trying to get help with the first question. It doesn't really make any sense.

OpenStudy (anonymous):

I taught as much too but the second one I got it right